# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ 2.3 Hyperparameters
Until now, we have assumed that the learning pipeline $\mathcal{I}$ is fixed. However, individual components of $\mathcal{I}$ usually have several hyperparameters (HPs), which determine their specific configuration and thus substantially influence the resulting prediction model. This also applies to the learning pipeline example considered in the previous section, for which possible HPs are shown in the left panel of Figure 1 (see below for further explanation). In contrast to the parameters $\boldsymbol{\theta}$, which are estimated as outputs of the learning pipeline, the HPs serve as inputs. This means that they must be specified before the learning pipeline is applied to the data set (Bischl et al.,, 2023).
@@ 2.3.1 Additional notation for HPs
The following notation is based on Feurer and Hutter, (2019). We denote the $j$th HP of a learning pipeline as $\lambda_{j}$, which is selected from its domain $\Lambda_{j}$ (i.e. $\lambda_{j}\in\Lambda_{j}$). The domain of $\lambda_{j}$ can generally be real-valued, integer-valued, binary, or categorical, as we will see in the examples given below. All $J$ HPs of a learning pipeline can be summarized as a vector $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{J})$ and their overall configuration space as $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\dots\times\Lambda_{J}$ (with $\boldsymbol{\lambda}\in\boldsymbol{\Lambda}$). Note that $\boldsymbol{\Lambda}$ may contain conditionality, meaning that some HPs might only be relevant when one or more other HPs are set to a certain value (see below for examples).
@@ 2.3.2 Algorithm HPs
Each learning algorithm usually has several HPs, which are specified by the software package used and can have a large impact on its complexity, speed, and other important properties of the algorithm (Bischl et al.,, 2023). For example, the HPs of the CART algorithm include the minimum number of observations in any terminal node ($\lambda_{minbucket})$, the maximum tree depth, with the root node counted as depth 0 ($\lambda_{maxdepth}$), and the factor by which a split needs to decrease the overall lack of fit to be attempted ($\lambda_{cp}$) (Therneau and Atkinson,, 2022). In the CART implementation of the R package mlr3 (Lang et al.,, 2019), the respective HP domains are $\Lambda_{minbucket}=\{1,\ldots,n_{\text{train}}\}$, $\Lambda_{maxdepth}=\{1,\ldots,30\}$ (both being integer-valued domains), and $\Lambda_{cp}=[0,1]$ (real-valued domain). Most algorithm HPs have default values that are specified by the software in which they are implemented (e.g., in mlr3, $\lambda_{minbucket}=7$ per default).
Note that since there is usually more than one algorithm suitable for a given prediction problem, the choice of algorithm can also be seen as an HP of the learning pipeline (with the HPs associated with each algorithm representing conditional HPs that are only relevant when the respective algorithm is used; Bischl et al.,, 2023). This creates an even more flexible but also complex learning pipeline, which is why, in this paper, we assume that the algorithm has already been selected.
@@ 2.3.3 Preprocessing HPs
As mentioned above, it is not only possible to specify learning algorithm HPs, but also preprocessing HPs (Binder and Pfisterer,, 2024; Bischl et al.,, 2023). In principle, whenever multiple options exist for performing a preprocessing step, these options can be considered as different HP values of the respective preprocessing step.
First, the choice of whether a preprocessing step $PS$ is applied at all can be considered as a binary HP $\lambda_{PS}$ with $\Lambda_{PS}=\{\mathrm{yes,no}\}$ (e.g., whether features should be log-transformed or not). Second, there is often more than one possible option for performing a preprocessing step. For example, the influence of outliers in features can be reduced by replacing all values that are outside the range $[x_{min},x_{max}]$ by $x_{min}$ and $x_{max}$, respectively (“winsorizing”; Steyerberg,, 2019). There are different options to specify $x_{min}$ and $x_{max}$, which means that $\lambda_{x_{min}}$ and $\lambda_{x_{max}}$ are HPs of the winsorizing preprocessing step (e.g., Steyerberg,, 2019 suggests percentiles such as $\lambda_{x_{min}}=1\text{st percentile}$ and $\lambda_{x_{max}}=99\text{th percentile}$).
[…]
@@
In contrast to algorithm HPs, it seems that preprocessing HPs – apart from those related to feature selection – are rarely discussed or referred to as such in ML applications (see, e.g., the systematic reviews of Dhiman et al., 2022a, and Andaur Navarro et al.,, 2023, where such terms were not mentioned). ML methods research usually also focuses on algorithm HPs rather than preprocessing HPs. An exception is the benchmark study by Stüber et al., (2023), which, among other factors, examines the impact of using principal component analysis in radiomics-based survival analysis.
@@ 2.3.4 Selection of HPs
While it is usually possible to leave all HPs at their respective default value, it is common to modify them in an attempt to optimize the prediction model generated by the learning pipeline. This can also be necessary if there is no specified default value. The term “optimization” here often refers to the predictive performance of the model but can also take into account other criteria such as simplicity, interpretability, or runtime to generate the model (Domingos,, 2012; Pfob et al.,, 2022; de Hond et al.,, 2022; Bischl et al.,, 2023). Note that the selection of HPs can be considered a “researcher degree of freedom” (Simmons et al.,, 2011), as it is one of many choices that users must make throughout the model development process (other choices are, e.g., how predictive performance is assessed; Hofman et al.,, 2017; Klau et al.,, 2019; Hosseini et al.,, 2020).
@@
We can distinguish between two primary types of HP selection: data-independent and data-dependent procedures. Data-independent HP selection does not make use of the data set $\mathcal{D}_{\text{train}}$ and is ideally based on the user’s knowledge about the data set and learning algorithm. For example, sensible algorithm HPs can be selected when users are experienced with the learning algorithm or when corresponding recommendations from the literature (e.g., previous benchmark studies) are available (Bartz et al.,, 2023; Bischl et al.,, 2023). Similarly, some preprocessing HPs may be inferred from substantive knowledge about the data set (e.g., which set of features should be selected) or knowledge about how the learning algorithm is affected by certain data set characteristics (e.g., whether the algorithm is sensitive to outliers in features, which requires some form of transformation; Kuhn and Johnson,, 2013). An example of data-independent HP selection on the basis of model simplicity is the specification of the maximum tree depth in the real-world prediction problem considered in Section 5, where the project team set the HP to $\lambda_{maxdepth}=4$ to ensure that the resulting decision tree can be implemented in clinical practice.
@@
In cases where users have insufficient knowledge about the data and learning algorithm to ensure a reasonable HP selection, but wish to avoid arbitrary or default HP values, it is possible to use the data set $\mathcal{D}_{\text{train}}$ to select optimal HP values. This process corresponds to a data-dependent HP selection, but terms such as HP tuning and (data-driven) HP optimization are more common (e.g., Bischl et al.,, 2023; Probst et al.,, 2019; Bartz et al.,, 2023). We will accordingly use the term HP tuning in the remainder of this paper. Note that HP tuning implies that not only the parameters $\boldsymbol{\theta}$ are estimated from the data set $\mathcal{D}_{\text{train}}$, but also one or more HPs in $\boldsymbol{\lambda}$. HP tuning thus generally complicates model generation and evaluation, which will be described in more detail in Section 4.
@@
Importantly, there are HPs that should not be selected through tuning. For learning algorithms, this includes, for example, the number of trees ($\lambda_{num.trees}$) in the random forest algorithm for classification problems: Due to the monotonous relation between $\lambda_{num.trees}$ and model performance in most cases, the largest computationally feasible number of trees should be chosen (Probst and Boulesteix,, 2018). Regarding preprocessing HPs, this typically applies to those associated with steps that alter the scope or interpretation of the prediction model (see Section 2.2.2). As such steps require careful specification, the corresponding HPs should be set based on user expertise (i.e. data-independently) rather than determined through tuning.
@@ 2.4.2 Model evaluation
Once the final prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ has been generated, the next important step is its evaluation. Since many algorithms yield black-box models that cannot be easily interpreted – and are thus difficult to assess for plausibility without additional tools (see, e.g., Molnar,, 2022) – a key quantity in the evaluation of a model is its prediction error. In the context of this work, we will accordingly use the term “model evaluation” synonymously with determining a model’s prediction error. The prediction error indicates how well a model performs on new observations that are independently drawn from the same distribution as the observations in $\mathcal{D}_{\text{train}}$ (i.e. from $\mathds{P}_{xy}$). It is specified with respect to a loss function $L$, which assesses the discrepancy between true outcomes and predictions and constitutes the performance measure. Formally, the prediction error of $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ can be defined as
@@
(Hastie et al.,, 2009; Boulesteix et al.,, 2015; Bischl et al.,, 2023). The loss function $L$ can be chosen according to the prediction problem being addressed. For instance, a common choice for $L$ in regression problems is the squared loss. In this case, the prediction error reflects the well-known mean squared error (MSE). Note that in equation (2), we assume for simplicity that $L$ corresponds to a point-wise loss function, although many commonly used performance measures (e.g., the area under the receiver operating characteristic curve, AUC) would necessitate a more general definition (provided in Bischl et al.,, 2023). Nonetheless, all following statements regarding the prediction error hold regardless of this simplified (and more common) representation.
An estimate of the prediction error in equation (2) can be obtained by using $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ to make predictions for an additional data set with new observations drawn from $\mathds{P}_{xy}$ (referred to as test data set $\mathcal{D}_{\text{test}}$). The prediction error can then be estimated by evaluating the loss function $L$ for each observation and calculating the average across all observations (again, assuming a point-wise loss; Bischl et al.,, 2023; Hastie et al.,, 2009). The resulting prediction error estimate for $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ can be denoted as $\widehat{\mathrm{PE}}(\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}},\mathcal{D}_{\text{test}})$. Note that the outcome values for $\mathcal{D}_{\text{test}}$ must be observed; otherwise, the loss function $L$ cannot be evaluated.
@@
Depending on the chosen evaluation procedure, a potential issue can be data leakage, which occurs whenever information about the designated $\mathcal{D}_{\text{test}}$ is improperly available during the generation of the model to be evaluated (Kapoor and Narayanan,, 2023; Kapoor et al.,, 2024; Kaufman et al.,, 2012; Rosenblatt et al.,, 2024; Hornung et al.,, 2023). Since in this case, the observations in $\mathcal{D}_{\text{test}}$ no longer truly represent new observations to which the model will be applied and the model thus has an unfair advantage when predicting these observations, the resulting prediction error estimate can be optimistically biased. Kapoor and Narayanan, (2023) identify three general types of data leakage, which may arise from: (i) overlap between the data used for model generation and evaluation, (ii) violation of the assumption that all observations are independently drawn from the same distribution, or (iii) use of illegitimate features. In this paper, we will focus on overlap-induced data leakage but provide additional information on the other two types in Supplementary Section A. Furthermore, we encounter an example of one of the other types in our empirical illustration in Section 5.
Finally, note that in some applications of ML (e.g., in the context of healthcare research), the process of assessing a model’s performance on observations from $\mathds{P}_{xy}$ is referred to as internal validation. This is in contrast to external validation, which evaluates how well the model predicts observations from different distributions (e.g., different time points or healthcare settings; de Hond et al.,, 2022; Collins et al., 2024a, ; Van Royen et al.,, 2023; Van Calster et al.,, 2023). As external validation is recommended to be performed in subsequent research only after successful internal validation (Collins et al., 2024a, ), we will focus on internal validation in this paper. Note that, in general, the term “evaluation” should be preferred over “validation” as the latter suggests that a “validated model” has a low prediction error, which is not necessarily the case (Collins et al., 2024a, ).
@@ 3.1 Model generation
The parameters of the learning algorithm are usually estimated based on a loss function $l$ that measures the discrepancy between the true outcome and a prediction vector for each observation $i$, i.e. $l(y^{(i)},f(\boldsymbol{x}^{(i)}))$. The algorithm parameters are then found by minimizing $\sum_{i=1}^{n_{\text{train}}}l(y^{(i)},f(\boldsymbol{x}^{(i)}))$ (see, e.g, Bischl et al.,, 2023 or Bartz et al.,, 2023 for more details). For example, in a regression problem where the learning algorithm corresponds to the CART algorithm, the splitting rules are found by minimizing the sum of squared errors and the prediction value for each terminal node corresponds to the mean of all outcome values in the respective node (Breiman et al.,, 1984). Note that the loss function $l$ may, but does not necessarily have to, align with the loss function $L$ from Section 2.4.2, which is used to estimate the prediction error.
When estimating the parameters, the learning pipeline may not only capture the signal in $\mathcal{D}_{\text{train}}$ which represents the true underlying data-generating mechanism $\mathds{P}_{xy}$, but it may also erroneously learn the specific pattern of noise (i.e. unexplained variation) in $\mathcal{D}_{\text{train}}$. The resulting prediction model is too adapted to $\mathcal{D}_{\text{train}}$ and will perform worse on new observations (drawn from $\mathds{P}_{xy}$) than on the observations in $\mathcal{D}_{\text{train}}$. This is a well-known problem in prediction model training and is commonly referred to as overfitting (e.g., Hastie et al.,, 2009; Kuhn and Johnson,, 2013; Steyerberg,, 2019; Bischl et al.,, 2023; Poldrack et al.,, 2020; de Hond et al.,, 2022). The risk of obtaining an overfitted prediction model depends on both the data set $\mathcal{D}_{\text{train}}$ (specifically on its signal-to-noise ratio, which tends to decrease as the number of observations decreases), and on the learning pipeline $\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{I}}}$ used to train the model (Poldrack et al.,, 2020; Lones,, 2024). The association between the characteristics of a learning pipeline and its tendency to overfit is not straightforward, but is related to factors such as the size of its hypothesis space (i.e. the number of prediction models that can be trained by $\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{I}}}$) and the procedure by which the model is chosen from the hypothesis space (e.g., whether the hypothesis space is searched exhaustively; Domingos,, 2012). These factors can vary greatly between learning pipelines, especially depending on the type of learning algorithm and the chosen HP values. Note that the learning pipeline may also suffer from underfitting rather than overfitting, which occurs if it is not flexible enough to adequately model the underlying data-generating mechanism (Hastie et al.,, 2009).

### v2

@@ 2.3 Hyperparameters
Until now, we have assumed that the learning pipeline $\mathcal{I}$ is fixed. However, individual components of $\mathcal{I}$ usually have several hyperparameters (HPs), which determine their specific configuration and thus substantially influence the resulting prediction model. This also applies to the learning pipeline example considered in the previous section, for which possible HPs are shown in the left panel of Figure 1 (see below for further explanation). In contrast to the parameters $\boldsymbol{\theta}$, which are estimated as outputs of the learning pipeline, the HPs serve as inputs. This means that they must be specified before the learning pipeline is applied to the data set [78].
@@ 2.3.1 Additional notation for HPs
The following notation is based on [95]. We denote the $j$th HP of a learning pipeline as $\lambda_{j}$, which is selected from its domain $\Lambda_{j}$ (i.e. $\lambda_{j}\in\Lambda_{j}$). The domain of $\lambda_{j}$ can generally be real-valued, integer-valued, binary, or categorical, as we will see in the examples given below. All $J$ HPs of a learning pipeline can be summarized as a vector $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{J})$ and their overall configuration space as $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\dots\times\Lambda_{J}$ (with $\boldsymbol{\lambda}\in\boldsymbol{\Lambda}$). Note that $\boldsymbol{\Lambda}$ may contain conditionality, meaning that some HPs might only be relevant when one or more other HPs are set to a certain value (see below for examples).
@@ 2.3.2 Algorithm HPs
Each learning algorithm usually has several HPs, which are specified by the software package used and can have a large impact on its complexity, speed, and other important properties of the algorithm [78]. For example, the HPs of the CART algorithm include the minimum number of observations in any terminal node ($\lambda_{minbucket})$, the maximum tree depth, with the root node counted as depth 0 ($\lambda_{maxdepth}$), and the factor by which a split needs to decrease the overall lack of fit to be attempted ($\lambda_{cp}$) [134]. In the CART implementation of the R package mlr3 [113], the respective HP domains are $\Lambda_{minbucket}=\{1,\ldots,n_{\text{train}}\}$, $\Lambda_{maxdepth}=\{1,\ldots,30\}$ (both being integer-valued domains), and $\Lambda_{cp}=[0,1]$ (real-valued domain). Most algorithm HPs have default values that are specified by the software in which they are implemented (e.g., in mlr3, $\lambda_{minbucket}=7$ per default).
Note that since there is usually more than one algorithm suitable for a given prediction problem, the choice of algorithm can also be seen as an HP of the learning pipeline (with the HPs associated with each algorithm representing conditional HPs that are only relevant when the respective algorithm is used; [78]). This creates an even more flexible but also complex learning pipeline, which is why, in this paper, we assume that the algorithm has already been selected.
@@ 2.3.3 Preprocessing HPs
As mentioned above, it is not only possible to specify learning algorithm HPs but also preprocessing HPs [77, 78]. In principle, whenever multiple options exist for performing a preprocessing step, these options can be considered as different HP values of the respective preprocessing step.
First, the choice of whether a preprocessing step $PS$ is applied at all can be considered as a binary HP $\lambda_{PS}$ with $\Lambda_{PS}=\{\mathrm{yes,no}\}$ (e.g., whether features should be log-transformed or not). Second, there is often more than one possible option for performing a preprocessing step. For example, the influence of outliers in features can be reduced by replacing all values that are outside the range $x_{min},x_{max}]$ by $x_{min}$ and $x_{max}$, respectively (“winsorizing”; [132]). There are different options to specify $x_{min}$ and $x_{max}$, which means that $\lambda_{x_{min}}$ and $\lambda_{x_{max}}$ are HPs of the winsorizing preprocessing step (e.g., [132], suggests percentiles such as $\lambda_{x_{min}}=1\text{st percentile}$ and $\lambda_{x_{max}}=99\text{th percentile}$).
Several possible options also exist for the imputation of missing feature values. For example, imputation can be based on the feature’s mean or median, or on a sampled value from its empirical distribution [135, as illustrated in]. This constitutes a (categorical) preprocessing HP $\lambda_{impute}$ with $\Lambda_{impute}=\{\text{mean, median, sample},\ldots\}$.
Another typical example of a preprocessing step with many possible options is feature selection. To define HPs in this context, we have to differentiate between filter and wrapper methods (the following explanations are based on [141], who also provides more details and additional examples). Filter methods are preprocessing steps that assign a numeric score to each feature (e.g., the correlation coefficient $\rho$ between each feature and the outcome) and select a set of features according to this score (e.g., all features with $\rho>0.2$). Consequently, the set of selected features is the parameter of the filter (i.e. $\theta_{filter}$, with, e.g., $\hat{\theta}_{filter}=\{x_{6},x_{8},x_{21},x_{25}\}$), while its specific configuration can be modified by its HPs. For example, there are different options to define the score ($\lambda_{filter_{1}}$, with $\Lambda_{filter_{1}}=\{\text{correlation, variance, importance score},\ldots\}$) and to select the features based on their score ($\lambda_{filter_{2}}$, with $\Lambda_{filter_{2}}=\{\text{top $r$ features, all features with a score}\geq\tau,\dots\}$, where $r$ and $\tau$ themselves are HPs that are conditional on $\lambda_{filter_{2}}$). Instead of using filter methods, it is also possible to directly specify the set of features that should be selected. In this case, the set of selected features is an input rather than an output of the learning pipeline and is therefore the HP ($\lambda_{features}$) of the feature selection step. For example, if only the features $x_{6},x_{9}$, and $x_{21}$ should be used by the learning algorithm, then $\lambda_{features}=\{x_{6},x_{9},x_{21}\}$. In many applications, $\lambda_{features}$ is not specified once by the user, but different values of $\lambda_{features}$ are tried and evaluated on $\mathcal{D}_{\text{train}}$. This process is referred to as a wrapper method but is, in fact, a special case of HP tuning, which will be discussed in Section 4.1.
@@
In contrast to algorithm HPs, it seems that preprocessing HPs—apart from those related to feature selection—are rarely discussed or referred to as such in ML applications (see, e.g., the systematic reviews of [88], and [74], where such terms were not mentioned). ML methods research usually also focuses on algorithm HPs rather than preprocessing HPs. An exception is the benchmark study by [133], which, among other factors, examines the impact of using principal component analysis in radiomics-based survival analysis.
@@ 2.3.4 Selection of HPs
While it is usually possible to leave all HPs at their respective default value, it is common to modify them in an attempt to optimize the prediction model generated by the learning pipeline. This can also be necessary if there is no specified default value. The term “optimization” here often refers to the predictive performance of the model but can also take into account other criteria such as simplicity, interpretability, or runtime to generate the model [90, 121, 86, 78]. Note that the selection of HPs can be considered a “researcher degree of freedom” [129], as it is one of many choices that users must make throughout the model development process (other choices are, e.g., how predictive performance is assessed; [101, 110, 104]).
@@
We can distinguish between two primary types of HP selection: data-independent and data-dependent procedures. Data-independent HP selection does not make use of the data set $\mathcal{D}_{\text{train}}$ and is ideally based on the user’s knowledge about the data set and learning algorithm. For example, sensible algorithm HPs can be selected when users are experienced with the learning algorithm or when corresponding recommendations from the literature (e.g., previous benchmark studies) are available [76, 78]. Similarly, some preprocessing HPs may be inferred from substantive knowledge about the data set (e.g., which set of features should be selected) or knowledge about how the learning algorithm is affected by certain data set characteristics (e.g., whether the algorithm is sensitive to outliers in features, which requires some form of transformation; [111]). An example of data-independent HP selection on the basis of model simplicity is the specification of the maximum tree depth in the real-world prediction problem considered in Section 5, where the project team set the HP to $\lambda_{maxdepth}=4$ to ensure that the resulting decision tree can be implemented in clinical practice.
@@
In cases where users have insufficient knowledge about the data and learning algorithm to ensure a reasonable HP selection but wish to avoid arbitrary or default HP values, it is possible to use the data set $\mathcal{D}_{\text{train}}$ to select optimal HP values. This process corresponds to a data-dependent HP selection, but terms such as HP tuning and (data-driven) HP optimization are more common [78, 124, 76, e.g.,]. We will accordingly use the term HP tuning in the remainder of this paper. Note that HP tuning implies that not only the parameters $\boldsymbol{\theta}$ are estimated from the data set $\mathcal{D}_{\text{train}}$ but also one or more HPs in $\boldsymbol{\lambda}$. HP tuning thus generally complicates model generation and evaluation, which will be described in more detail in Section 4.
@@
Importantly, there are HPs that should not be selected through tuning. For learning algorithms, this includes, for example, the number of trees ($\lambda_{num.trees}$) in the random forest algorithm for classification problems: Due to the monotonous relation between $\lambda_{num.trees}$ and model performance in most cases, the largest computationally feasible number of trees should be chosen [123]. Regarding preprocessing HPs, this typically applies to those associated with steps that alter the scope or interpretation of the prediction model (see Section 2.2.2). As such steps require careful specification, the corresponding HPs should be set based on user expertise (i.e. data-independently) rather than determined through tuning.
@@ 2.4.2 Model evaluation
Once the final prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ has been generated, the next important step is its evaluation. Since many algorithms yield black-box models that cannot be easily interpreted, and are thus difficult to assess for plausibility without additional tools (see, e.g., [115]), a key quantity in the evaluation of a model is its prediction error. In the context of this work, we will accordingly use the term “model evaluation” synonymously with determining a model’s prediction error. The prediction error indicates how well a model performs on new observations that are independently drawn from the same distribution as the observations in $\mathcal{D}_{\text{train}}$ (i.e. from $\mathds{P}_{xy}$). It is specified with respect to a loss function $L$, which assesses the discrepancy between true outcomes and predictions and constitutes the performance measure. Formally, the prediction error of $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ can be defined as
@@
[98, 79, 78]. The loss function $L$ can be chosen according to the prediction problem being addressed. For instance, a common choice for $L$ in regression problems is the squared loss. In this case, the prediction error reflects the well-known mean squared error (MSE). Note that in equation (2), we assume for simplicity that $L$ corresponds to a point-wise loss function, although many commonly used performance measures (e.g., the area under the receiver operating characteristic curve, AUC) would necessitate a more general definition (provided in [78]). Nonetheless, all following statements regarding the prediction error hold regardless of this simplified (and more common) representation.
An estimate of the prediction error in equation (2) can be obtained by using $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ to make predictions for an additional data set with new observations drawn from $\mathds{P}_{xy}$ (referred to as test data set $\mathcal{D}_{\text{test}}$). The prediction error can then be estimated by evaluating the loss function $L$ for each observation and calculating the average across all observations (again, assuming a point-wise loss; [78, 98]). The resulting prediction error estimate for $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ can be denoted as $\widehat{\mathrm{PE}}(\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}},\mathcal{D}_{\text{test}})$. Note that the outcome values for $\mathcal{D}_{\text{test}}$ must be observed; otherwise, the loss function $L$ cannot be evaluated.
@@
Depending on the chosen evaluation procedure, a potential issue can be data leakage, which occurs whenever information about the designated $\mathcal{D}_{\text{test}}$ is improperly available during the generation of the model to be evaluated [108, 107, 109, 127, 103]. Since, in this case, the observations in $\mathcal{D}_{\text{test}}$ no longer truly represent new observations to which the model will be applied, and the model thus has an unfair advantage when predicting these observations, the resulting prediction error estimate can be optimistically biased. [108] identify three general types of data leakage, which may arise from: (i) overlap between the data used for model generation and evaluation, (ii) violation of the assumption that all observations are independently drawn from the same distribution, or (iii) use of illegitimate features. In this paper, we will focus on overlap-induced data leakage but provide additional information on the other two types in Supplementary Section A. Furthermore, we encounter an example of one of the other types in our empirical illustration in Section 5.
Finally, note that in some applications of ML (e.g., in the context of healthcare research), the process of assessing a model’s performance on observations from $\mathds{P}_{xy}$ is referred to as internal validation. This is in contrast to external validation, which evaluates how well the model predicts observations from different distributions (e.g., different time points or healthcare settings; [86, 84, 137, 136]). As external validation is recommended to be performed in subsequent research only after successful internal validation [84], we will focus on internal validation in this paper. Note that, in general, the term “evaluation” should be preferred over “validation” as the latter suggests that a “validated model” has a low prediction error, which is not necessarily the case [84].
@@ 3.1 Model generation
The parameters of the learning algorithm are usually estimated based on a loss function $l$ that measures the discrepancy between the true outcome and a prediction vector for each observation $i$, i.e. $l(y^{(i)},f(\boldsymbol{x}^{(i)}))$. The algorithm parameters are then found by minimizing $\sum_{i=1}^{n_{\text{train}}}l(y^{(i)},f(\boldsymbol{x}^{(i)}))$ (see, e.g., [78], or [76], for more details). For example, in a regression problem where the learning algorithm corresponds to the CART algorithm, the splitting rules are found by minimizing the sum of squared errors and the prediction value for each terminal node corresponds to the mean of all outcome values in the respective node [81]. Note that the loss function $l$ may, but does not necessarily have to, align with the loss function $L$ from Section 2.4.2, which is used to estimate the prediction error.
When estimating the parameters, the learning pipeline may not only capture the signal in $\mathcal{D}_{\text{train}}$ which represents the true underlying data-generating mechanism $\mathds{P}_{xy}$, but it may also erroneously learn the specific pattern of noise (i.e. unexplained variation) in $\mathcal{D}_{\text{train}}$. The resulting prediction model is too adapted to $\mathcal{D}_{\text{train}}$ and will perform worse on new observations (drawn from $\mathds{P}_{xy}$) than on the observations in $\mathcal{D}_{\text{train}}$. This is a well-known problem in prediction model training and is commonly referred to as overfitting [98, 111, 132, 78, 122, 86, e.g.,]. The risk of obtaining an overfitted prediction model depends on both the data set $\mathcal{D}_{\text{train}}$ (specifically on its signal-to-noise ratio, which tends to decrease as the number of observations decreases) and on the learning pipeline $\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{I}}}$ used to train the model [122, 114]. The association between the characteristics of a learning pipeline and its tendency to overfit is not straightforward, but it is related to factors such as the size of its hypothesis space (i.e. the number of prediction models that can be trained by $\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{I}}}$) and the procedure by which the model is chosen from the hypothesis space (e.g., whether the hypothesis space is searched exhaustively; [90]). These factors can vary greatly between learning pipelines, especially depending on the type of learning algorithm and the chosen HP values. Note that the learning pipeline may also suffer from underfitting rather than overfitting, which occurs if it is not flexible enough to adequately model the underlying data-generating mechanism [98].